OpenStudy (anonymous):

f(x) = 19xe^(1 − x^2)

OpenStudy (campbell_st):

what do you need to do...?

OpenStudy (anonymous):

Find the exact location of all the relative and absolute extrema of the function.

OpenStudy (campbell_st):

well perhaps you should graph it 1st here is a sight I use https://www.desmos.com/calculator then you can see them... makes the calculus a little easier

OpenStudy (anonymous):

i did that and i typed in what i got and it came out wrong>.<

OpenStudy (campbell_st):

type in y = 19xe^(1 - x^2) and it should work

OpenStudy (anonymous):

the graphing works! but the point are wrong when i put them as my answer!

OpenStudy (campbell_st):

is there a domain in your question..?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

f(x) = x4 − 8x3 with domain [−1, +∞) do you know how to do this one?

OpenStudy (anonymous):

f(x)=x^4 - 8x^3 domain -1,+infinity

OpenStudy (aum):

f(x) = x^4 - 8x^3 f'(x) = 4x^3 - 24x^2 = 4x^2(x - 6) = 0 x = 0, x = 6 are critical points. f''(x) = 12x^2 - 48x f''(0) = 0 (therefore x = 0 is not a local maxima / minima_ f''(6) = 12*36 - 48*6 = 144. f''(6) > 0, therefore x = 6 is a local minima.

OpenStudy (anonymous):

so what is the y?

OpenStudy (anonymous):

im so lost and my homework is due at 10 am:/

OpenStudy (aum):

at x = 6, f(6) = 6^4 - 8*6^3 = -432.

OpenStudy (aum):

So when x = 6, the corresponding y value is -432.

OpenStudy (anonymous):

so its telling me that its wrong>.<

OpenStudy (anonymous):

wait just kidding it took it!

OpenStudy (anonymous):

f(x) = x ln x with domain (0, +∞) do you know how to do this one by chance?

OpenStudy (anonymous):

domain (0,+infinity)